OpenStudy (anonymous):

The surface area of a rectangular prism is found using the formula SA = 2(lw + lh + wh), where l is length, w is width, and h is height. What is the surface area (in square meters) of a rectangular prism with length 5 m, width 12 m, and height 6 m? A. 92 B. 102 C. 324 D. 360

OpenStudy (anonymous):

SA=2(lw+lh+wh) Which is, in this case SA=2(5∗12+5∗6+12∗6) Which is the same as SA=2(60+30+72) Which is SA=2(162)=324 Which is answer choice C.